OpenStudy (swaggyswags):

Consider a graph of the equation y = x + 6. What is the y-intercept? −1 1 −6 6

OpenStudy (princeharryyy):

put x =0 for y intercept

OpenStudy (princeharryyy):

y = 6

OpenStudy (swaggyswags):

Thanks, could you help me with a few more?

OpenStudy (princeharryyy):

okay

OpenStudy (swaggyswags):

What is the equation in point-slope form of the line passing through (2, 5) and (−1, 8)? y + 5 = 1(x + 2) y + 8 = −1(x − 1) y − 5 = −1(x − 2) y − 8 = 1(x − 2)

OpenStudy (princeharryyy):

C

OpenStudy (swaggyswags):

How did you get that?

OpenStudy (princeharryyy):